OpenStudy (anonymous):

A person can see the top of a building at an angle of 65 degrees. The person is standing 50 ft away from the building and has an eye level of 5 ft. How tall is the building to the nearest tenth of a foot?

OpenStudy (anonymous):

@e.mccormick Would you mind helping me with this and four other questions if you can?

OpenStudy (johnweldon1993):

 So you have an angle...a side adjacent to that angle...and you want the side opposite that angle.. What trigonometric function uses opposite and adjacent?

OpenStudy (anonymous):

Uh...Sin Cos Tan???

OpenStudy (johnweldon1993):

Hint* Remember SOH CAH TOA it helps remember the trig functions S sin = O opposite H hypotenuse C cos = A adjacent H hypotenuse T tan = O opposite A adjacent So which one is it?

OpenStudy (anonymous):

I'm not sure, to be honest. I'm really bad at math.

OpenStudy (anonymous):

@johnweldon1993 Uhm, Are you done helping or??

OpenStudy (hoblos):

tan = opposite/adjacent so tan65 = h/50 h=?

OpenStudy (anonymous):

Uhm...I don't know how to find that answer.

OpenStudy (hoblos):

you have tan65 = h/50 can you find h from that ?

OpenStudy (anonymous):

I don't know how.

OpenStudy (hoblos):

multiply both sides by 50 what do you get ?

OpenStudy (anonymous):

Oh! 3250, 2500. Sorry for my slowness, I'm lagging and working on biology.

OpenStudy (anonymous):

@Opcode Could you please help?

OpenStudy (anonymous):

@whpalmer4 Ahhh..Can you help? x-x....